OpenStudy (anonymous):

if 5^-x=3 what does 5^3x equal?

OpenStudy (johnweldon1993):

So if \[\large 5^{-x} = \frac{1}{5^x} = 3\] We would solve for 'x' and use that in the new equation we have what does 'x' equal there?

OpenStudy (cruffo):

If \(5^{-x} = 3\) then \(5^x = \dfrac{1}{3}\). Then cube both sides of the equation to get \[5^{3x} = \left(\dfrac{1}{3}\right)^3 = \dfrac{1}{27}\]

OpenStudy (anonymous):

Ohh okay thanks